OpenStudy (anonymous):

how to calculate in comp science im not aable to solve this 5+4/3*(5-3)/2

OpenStudy (saifoo.khan):

\[5 + \frac43 \times \frac{(5-3)}{2}\]

OpenStudy (saifoo.khan):

Use PEDMAS.

OpenStudy (anonymous):

simple !! @saifoo.khan take charge!! its u in the show 2 help

OpenStudy (saifoo.khan):

Parenthesis, Exponents, Division, Multiplication, Add Subtract.

OpenStudy (anonymous):

answer is ???

OpenStudy (saifoo.khan):

Calculate it to find out.

OpenStudy (anonymous):

answer is: try it

OpenStudy (anonymous):

what do you mean by computer science?

OpenStudy (anonymous):

6 sahi hai GOOD tnx

OpenStudy (anonymous):

5/3 *6 = ? according to PEDMAS is it 9.6?? if not then please explain how to use PEDMAS I hv started C++ since 2 hr

OpenStudy (anonymous):

76.6227766

OpenStudy (anonymous):

is answer

OpenStudy (anonymous):

nooo answer is 6

OpenStudy (anonymous):

:-(

OpenStudy (lgbasallote):

i got 5.66666 when rounded off is 6 does that count? haha

OpenStudy (anonymous):

dnt know, but the answer for 6.0 * 5/3 is 10.0 and 6 * 5/3 is 10 AND IM CONFUSED at the end!!

OpenStudy (lgbasallote):

6* 5/3 = (6)(5)/3 = 30/3 = 10 that i know

OpenStudy (lalaly):

5-3=2 3*2=6 4/6/2=4/3 5+4/3 = 6.33333

OpenStudy (lgbasallote):

but it says 6.0 must be the answer...

OpenStudy (lalaly):

6.3 in c++ when returned as an integer is 6

OpenStudy (anonymous):

then how 5/3 *6 = 6 , its given in answers

OpenStudy (lalaly):

i dont think thats right

OpenStudy (anonymous):

r u sure

OpenStudy (lalaly):

5/3*6 = 5/18 that is not 6

OpenStudy (anonymous):

also 6.0 * 5/3 = 10.0 n 6 * 5/3 = 10 how??

OpenStudy (lalaly):

thats why i said its wrong,,, the only way to do it is how i did it

OpenStudy (anonymous):

lol because 6*5/3 is 10

OpenStudy (anonymous):

what language do you code in?

OpenStudy (lalaly):

c++ he said

OpenStudy (anonymous):

hey hey when im SOLVING in WINDOWScalclater ans are correct and same as the book :-D :-D

OpenStudy (lalaly):

lol i dont know then,,, -.-

OpenStudy (anonymous):

Oh c++ hmm do you know the difference between int and float? when you add that decimal point the numbers gets stored as float.

OpenStudy (anonymous):

and int only account for the whole numbers.

OpenStudy (anonymous):

that point is clear( 10.0 and 10) FLOATING CONSTANT, we represent 10.0 and int constant 10 no fractional or decimals

OpenStudy (phi):

This may be confusion 5.0/3.0*6.0 but when you have operations at the same level (as are division and multiplication), you evaluate them left to right: e.g. \[ \frac{5}{3} \cdot 6 = 10 \] Here is C code to show the evaluation #include <stdio.h> main() { printf("Hello, World!\n"); printf("%s, %f\n","5.0/3.0*6.0", 5.0/3.0*6.0); } code$ cc hello.c -o hello -lm code$ hello Hello, World! 5.0/3.0*6.0, 10.000000

OpenStudy (anonymous):

i got it 5/3 is 1.6666666666667 so 1.666666667 *6 = 10 6.0 * 5/3 = 6.0 **1.66666667 = 10.0 BCZ FLOAT 6 * 5/3 = 10 simply bcz its in int and probably answer in the book is wrong for 5/3 *6 , answer given is 6 but it will be 10 now im sure

OpenStudy (anonymous):

DHANYVAAD EVERYONE tnx

OpenStudy (phi):

If you evaluate integers 5/3*6 then 5/3 -->1 (you lose the remainder) and 1*6=6 #include <stdio.h> main() { printf("Hello, World!\n"); printf("%s, %f\n","5.0/3.0*6.0", 5.0/3.0*6.0); printf("%s, %d\n","5/3*6", 5/3*6); } code$ hello Hello, World! 5.0/3.0*6.0, 10.000000 5/3*6, 6

OpenStudy (phi):

You say 6.0 * 5/3 is 10.0 <-- notice the 6.0, this causes the expression to be evaluated as a float 5/3 *6 = 6 <--- evaluated as integers

OpenStudy (anonymous):

yes i got it